OpenStudy (anonymous):

A graph has equation y=2x^2-4. What is the equation of the new graph when it is translated 2 units in the positive x direction

OpenStudy (anonymous):

2 units to the right is \[f(x-2)\] \[y=2(x-2)^{2}-4\]

OpenStudy (anonymous):

these can be opened to \[y=2x^2-8x+4\]

OpenStudy (anonymous):

these can be seen as the TP(0;-4) is now (2;-4)

OpenStudy (anonymous):

do any of u mind explaining?:D

OpenStudy (anonymous):

translating a graph can be done in two ways >shifting it along the y-axis which affects \[f(x)\] if a graph is moved 2 units up \[f(x)+2\] \[y=2x^2-4+2\] and two units down \[f(x)-2\] \[y=2x^2-4-2\] >but when pelleting along the x-axis it is always oppsite ie.2 units right(the positive x direction) affects x\[f(x-2)=2(x-2)^2-4\] ie.2 units left(the positive x direction) affects x\[f(x+2)=2(x+2)^2-4\]

OpenStudy (anonymous):

why is it f(x-2)?

OpenStudy (turingtest):

\[f(x-a)\]shifts the graph of \(f(x)\) by \(a\) units to the right why? consider some random graph...at x=a I have marked the point now what about the graph of \(f(x-a)\) ? well, if we change our graph from f(x) to f(x-a) then the point f(a) becomes f(a-a)=f(0) so the point that was at x=a is now f(0) instead of f(a) let's see what that looks like...